OpenStudy (zubhanwc3):

The grand canyon is 1600 meters deep at its deepest point. A rock is dropped from the rim above this point. Express the height of the rock as a function of the time t in seconds. How long will it take the rock to hit the canyon floor? a(t) = -9.8m/sec/sec due to gravity.

OpenStudy (anonymous):

I'm assuming you're not meant to derive the kinematics relationship \[\Delta x = \frac{ 1 }{ 2 }at^{2}\] so we're just gonna run with it. if we call x the displacement from the rim to falling rock, and plug in our acceleration, the equation for x vs. t is \[\ x = \frac{ -9.8 }{ 2 }t^{2}\] To solve for the time to the bottom, let x = -1600 \[\ -1600 = \frac{ -9.8 }{ 2 }at^{2}\] \[\ \sqrt \frac {-1600*2}{-9.8} = t\] t = 18 seconds

OpenStudy (zubhanwc3):

thankyou very much.